# The Coralscapes Dataset: Semantic Scene Understanding in Coral Reefs

### v1

@@ I Introduction
However, there are regions, species, and genotypes within species that are more heat-resistant, holding promise to withstand the projected ocean warming induced by climate change [13]. The reefs of the northern Red Sea, for example, are projected to resist up to 5°C of warming [29, 77, 45, 60, 71]. These heat-resistant reefs must be protected from local stressors to ensure the survival of coral reef ecosystems on the planet. To inform effective conservation or restoration strategies, reefs must be monitored at the highest spatial and temporal resolution. As remote sensing methods are limited in resolution, constrained to very shallow areas, and rely on good atmospheric, weather, and water conditions [49, 6, 18, 41], in situ surveying remains the de facto method for monitoring coral reefs.
Computer vision emerges as a promising tool to scale coral reef monitoring, through tackling the bottleneck imposed by the needed expert time for identifying corals and estimating their abundance on images, as well as the development of visual simultaneous localization and mapping (SLAM) systems that can be used on autonomous underwater vehicles (AUVs) and other robots. However, existing datasets in the coral reef domain are limited to narrow domains such as orthomosaics or photo quadrats [5, 27], to sparse point labels [8, 9] or image classification [68, 39], or in their size and diversity, inhibiting the widespread deployment of computer vision in coral reef science and conservation.
@@
  • The largest expert-annotated semantic segmentation dataset spanning 2075 images at 1024$\times$2048px resolution, containing 174k polygons over 39 classes, labeled in a consistent and speculation-free manner.
@@
II Related Work
II-A Large Datasets for Semantic Segmentation
Large benchmark datasets have played a key role in deep learning’s rapid advancement in semantic segmentation many other computer vision tasks. Early datasets such as PASCAL VOC [28] provided pixel-wise annotations for object categories, even before deep learning was ubiquitous. The introduction of Cityscapes [23], a dataset of 5000 images at 1024$\times$2048px with fine-grained semantic annotations in urban driving scenarios, set a precedent for large-scale semantic segmentation datasets. Notably, Cityscapes became the standard benchmark for evaluating segmentation networks, influencing the design of widely used neural network architectures such as the DeepLab [19] family of models [20, 21] and Pyramid Scene Parsing Networks [83].
II-B Coral Reef Classification & Segmentation
Dataset # Images Annotation Type # Annotations # Annotated Pixels # Classes Comment Spatial Split Openly Available
@@ II-B Coral Reef Classification & Segmentation
Benthoz-2015 [12] 9874 Sparse 407,968 - 148 Ortho-photos Yes Yes
@@
Table I: Comparison of existing datasets for computer vision in coral reefs in terms of annotation type, size, classes, availability of disjoint geographic locations allowing a spatially disjoint train/test split, and open availability.
In the context of coral reefs, existing expert-labeled datasets provide either only image-level labels [73, 39, 68, 31] on close-up images of corals, a few sparse labeled pixels on entire images [10, 43, 8, 9], or focus on segmentation in restricted domains such as photo quadrats or orthomosaics [82, 27]. Table I provides a comparison of existing datasets, including the BenthicNet [51] dataset (subselected for coral images, i.e. all sites with at least one hard coral label and shallower than 60m), and CoralSCOP [84], which provides a model similar to Segment Anything [44] trained on annotated masks of corals (but no other benthic classes).
Most existing datasets do not split the training, validation, and testing data according to different geographic sites, leading to a likely overestimation of the classification performance compared to real-world applications in unseen reefs. Furthermore, the larger, but sparsely annotated datasets [51, 12, 32] are essentially crowd-sourced from various annotation teams, leading to a large and noisy label set, that is challenging to transfer to new biogeographic regions. Lastly, except for the SUIM dataset [38], which provides dense segmentation labels focusing on robot-human interaction on diving missions and has no dedicated classes for live corals, no existing segmentation datasets are general-purpose in the sense of diverse camera angles, distances to the substrate, environmental conditions (water turbidity, color, lighting), and with all relevant classes visible on the images annotated.
@@
III Dataset
III-A Data Collection
All imagery was collected during scuba dives at 35 sites in coral reefs of Djibouti, Eritrea, Sudan, Jordan, and Israel using GoPro Hero 10 cameras. The annotated images are video frames originally taken at 1080 $\times$ 1920px resolution and 30 frames per second in the linear lens setting. The videos were taken in a diverse set of reefs: from thriving reefs with very high live coral cover to devastated environments with severe bleaching and coral mortality, from high visibility in good lighting conditions to turbid scenes in low-light settings. The videos were acquired during transect surveying campaigns of the Transnational Red Sea Center hosted at the École Polytechnique Fédérale de Lausanne (EPFL). Therefore, Coralscapes includes images including transect tools, transect lines, divers, and some human-made structures commonly encountered in reef monitoring settings. To protect the coral reefs in Coralscapes from overtourism or illegal poaching activities, and in accordance with local research permit authorities, the location of the sites is withheld and instead replaced by an ID.
III-B Annotation
@@
To align to all those requirements, the annotation was done in a speculation-free manner, enforcing a conservative semantic segmentation in terms of taxonomic depth. For example, when a branching coral was too far away from the camera to decide with certainty whether it is alive or dead, it was labeled as ‘background’. Viewed closer in another video frame, the same coral may be assigned to its growth form (branching, massive, etc.) and whether it is alive or dead. A precise genus label was assigned only when the genus was clearly distinguishable. In-depth details of the annotated classes is given in Appendix -B.
Figure 2: Histogram of the image area covered by annotations (left) and of the number of polygons present per image (right), highlighting the complexity of semantic segmentation in Coralscapes, compared to the CoralSCOP dataset [84].
@@
Annotation of the selected 39 classes with segmentation masks was done by four experts on coral reefs who were also present during data collection, and six additional trained annotators. The specific image frames for annotation were selected from over 200 hours of video material in order to capture the diversity of the scenes, while obtaining a good coverage of the 39 classes. The CVAT annotation interface [24] was used, which allowed annotators to interact with the Segment Anything Model [44]. While this helped to segment some objects with salient boundaries, it proved to be largely unhelpful for classes with less apparent boundaries and those appearing as small polygons, where polygons were drawn by hand. The annotation process emphasized high-quality annotations, meaning that there was a focus on annotating each frame with many polygons and on covering much of the image area, instead of annotating many images sparsely, as shown in Fig. 2. All polygons were visually verified to be correctly delineated and assigned the correct class.
In terms of class presence, the annotations of Coralscapes capture the strong imbalance in reef scenes: despite making an effort to prioritize the selection of frames with rare classes, the most common class has over 400 times more annotated polygons than the rarest class, as shown in Figure 1. Similarly, the granularity and size of the visible benthic classes varies dramatically: even though a class may appear very often, its total number of annotated pixels may still be low because it often appears in small polygons, and vice versa, as highlighted in Figure 3. For example, the median size of a ‘fish’ polygon is more than 120 times smaller than the median ‘background’ polygon. This imbalance in terms of representation and scale contributes to making Coralscapes a challenging benchmark for semantic segmentation. Additional statistics of Coralscapes (pixel counts, class frequency at image level) are provided in Appendix -A.
III-C Dataset Structure
@@
To facilitate the usability of Coralscapes, the structure of the dataset mimics that of the widely used Cityscapes benchmark: images are provided as 8-bit PNG images resized to 1024$\times$2048px resolution, and the 19 preceding and 10 trailing video frames are available for download (at 30 FPS). We make Coralscapes easily available: frames and their segmentation masks are available on Huggingface [^1] https://huggingface.co/datasets/EPFL-ECEO/coralscapes, from where they can be loaded to Python in one line, and on Zenodo [^2] https://zenodo.org/records/15061505, where we also provide the preceding and trailing frames, the frames at original 1080$\times$1920px resolution, and benchmarked model checkpoints.
IV Benchmarking
@@
Table II: Quantitative results on Coralscapes segmentation, best shown in bold, second best underlined.
@@
SegFormer MiT-B2
SegFormer MiT-B5
Figure 4: Qualitative samples from the Coralscapes test set. Additional samples provided in the Appendix.
@@
Table III: Transfer learning on UCSD Mosaics using 5 / 10 / 25 / 300 / all labeled pixels per image. Models marked with (Ours) are pre-trained on Coralscapes.
@@
Whenever available, we follow the original authors’ training procedures on Cityscapes in terms of hyperparameters, augmentations, and inference/evaluation strategy, unless otherwise specified. Implementation details for all benchmarked models are provided in Appendix A. To choose the number of training epochs, we train each model on the training set only and evaluate on the validation set. We then re-train each model on the train+validation sets with the number of epochs leading to the highest validation mIoU (smoothed across 10 evaluation epochs in order to reduce the effect of noise). Quantitative results are shown in Table II, where we report performance of the models averaged over last three evaluated epochs, in order to reduce noise without training an ensemble of classifiers or cherry-picking results. A qualitative comparison of models is shown in Fig 4.
@@
While the SegFormer models showed the best performance in terms of mIoU, the models with a DINOv2 backbone showed the best performance in terms of pixel accuracy. This likely stems from the fact that the DINOv2 models predict the segmentation maps for the image at once, where the global structure of the image helps determine the slightly subjective delineation of ‘background’, as opposed of the strided patch predictions from SegFormer, which processes the image at higher resolution and can predict the fine-grained classes better. We find that deviating from the training procedure of SegFormer on the Cityscapes dataset by increasing the strength of data augmentations, the performance can be further improved, yielding the best model.
V Transfer Learning & Applications
In this section, we evaluate the potential of transferring models trained on Coralscapes to three downstream tasks.
@@ V Transfer Learning & Applications
UCSD Mosaics Transfer Learning    We evaluate models trained on Coralscapes in a transfer learning setting on the UCSD Mosaics dataset [27], which consists of 16 orthomosaics densely annotated with 33 benthic classes and one class for the remaining substrate. We follow [5, 64, 65] in dividing the orthomosaics into patches of size 512$\times$512px, and remove patches with corrupted ground-truth masks [64], leaving 3974 training and 696 test patches. We evaluate both in a dense segmentation setting, as well as in a commonly performed sparse-to-dense setting, in which the training annotations are sparsified, but evaluation is performed on the unchanged dense labels. In existing works [5, 64, 65], this setting is tackled by a two-step procedure, where in a first step, algorithms propagate the sparse labels into denser masks, and in a second step, a segmentation neural network is trained on the obtained denser masks. Our approach is conceptually simpler: we train segmentation models directly on the sparsified labels, ignoring the unlabeled pixels. We select 5, 10, 25, and 300 pixel locations per orthomosaic patch uniformly at random, comparing models pre-trained on Coralscapes (with the last layer re-initialized) against off-the-shelf models with a pre-trained backbone. The results, shown in Table III, show that pre-training on Coralscapes substantially improves segmentation with DeepLabV3+, particularly in highly sparse label regimes (5, 10, 25). When fine-tuning all parameters of a SegFormer MiT-B2, pre-training on Coralscapes improves performance in the sparse regimes, but not in the less sparse or dense (“All” in Table III) settings. When training only the lightweight decoder on the MiT-B5 encoder weights obtained by pre-training on Coralscapes, we observe a consistent performance increase compared to an ImageNet pre-trained encoder. Interestingly, even an off-the-shelf DeepLabV3+ trained on sparse randomly located labels outperforms existing two-step procedures in sparse label regimes, despite [65] showing that selecting pixels on a grid or with a human in the loop instead of randomly is beneficial.
Crown of Thorns Starfish Survey    Acanthaster planci, colloquially known as the Crown of Thorns starfish (COTS), is a common reef inhabitant in the entire Indo-Pacific, where it preys on hard corals. While in normal population sizes, COTS are a healthy part of the ecosystem, there are outbreaks of COTS, in which their population increases dramatically and the hard coral cover declines, which can lead to devastating consequences for entire reefs [34, 63]. Interventions against COTS outbreaks are most effective in the early stages of an outbreak [62, 79]: computer vision can help to rapidly detect such outbreaks by automatically analyzing imagery from reefs for COTS abundance. In [50], a dataset of three videos from manta tows (a snorkeler with a camera, towed behind a boat) in Australia were annotated with bounding boxes of COTS. The videos contain 12347, 11374, and 10759 frames at 720 $\times$ 1280 px with 3065, 6384, and 2449 bounding-box annotations of COTS respectively. We use the bounding-boxes as segmentation labels to fine-tune a SegFormer-MiT-B5 pre-trained on Coralscapes for 10 epochs, with bounding boxes being drawn around connected components of the prediction for evaluation. The results (Table IV) show that pre-training on Coralscapes leads to a substantial improvement compared to using an ImageNet pre-trained backbone. Compared to a Yolo-V8-L baseline, the transfer-learned segmentation model consistently improves the mAP@50, but the mAP@50-95, which is more sensitive to the bounding box locations, is lower. This is likely due to the naive approach to transforming segmentations into bounding boxes.
@@
Table IV: COTS detection results in a leave-one-out test-set setting for each of the three videos.
Underwater 3D Mapping from Videos.    Even though 3D photogrammetry is widely used to map coral reefs [16, 46, 74, 67, 15], commonly used pipelines are designed for in-air mapping and often struggle in underwater environments. The detrimental effects of the water column include changes in the color of objects relative to the distance to the camera and induced blur [4, 70]. Dynamic objects such as moving fish or divers also lead to artifacts in the 3D reconstructions or even complete failure. Therefore, image collections are usually highly curated [7] to exclude images with such objects or with the water column visible.
@@
VI Conclusion and Future Work

### v2

@@ 1 Introduction
However, there are regions, species, and genotypes within species that are more heat-resistant, holding promise to withstand the projected ocean warming induced by climate change [13]. The reefs of the northern Red Sea, for example, are projected to resist up to 5°C of warming [30, 83, 48, 66, 77]. These heat-resistant reefs must be protected from local stressors to ensure the survival of coral reef ecosystems on the planet. To inform effective conservation or restoration strategies, reefs must be monitored at the highest spatial and temporal resolution. As aerial and satellite remote sensing methods are limited in resolution, constrained to very shallow areas, and rely on good atmospheric, weather, and water conditions [52, 5, 19, 44], in situ surveying remains the de facto method for monitoring coral reefs.
Computer vision emerges as a promising tool to scale coral reef monitoring, through tackling the bottleneck imposed by the needed expert time for identifying corals and estimating their abundance on images [8, 9], as well as the development of visual simultaneous localization and mapping (SLAM) systems [61, 62] that can be used on autonomous underwater vehicles (AUVs) and other robots [60, 33]. However, existing datasets in the coral reef domain are limited to narrow domains such as orthomosaics or photo quadrats [4, 28], limited to sparse point labels [8, 9] or image classification [74, 42], or in their size and diversity, inhibiting the widespread deployment of computer vision in coral reef science and conservation.
@@
[…]
@@
Dataset # Images Annotation Type # Annotations
# Annotated
Pixels
# Classes Comment
Spatial
Split
Openly
Available
@@
Benthoz-2015 [12] 9874 Sparse 407,968 - 147 Ortho-photos Yes Yes
@@
Table 1: Comparison of existing datasets for computer vision in coral reefs in terms of annotation type, size, classes, availability of disjoint geographic locations allowing a spatially disjoint train/test split, and open availability.
2 Related Work
2.1 Large Datasets for Semantic Segmentation
[…]
2.2 Coral Reef Classification & Segmentation
In the context of coral reefs, existing expert-labeled datasets provide either only image-level labels [79, 42, 74, 32] on close-up images of corals, a few sparse labeled pixels on entire images [10, 46, 8, 9], or focus on segmentation in restricted domains such as photo quadrats or orthomosaics [89, 28]. Table 1 provides a comparison of existing datasets, including the BenthicNet [54] dataset (subselected for coral images, i.e. all sites with at least one hard coral label and shallower than 60m), and CoralSCOP [91], which provides a model similar to Segment Anything [47] trained on annotated masks of corals (but no other benthic classes).
Most existing datasets do not split the training, validation, and testing data according to different geographic sites, leading to a likely overestimation of the classification performance compared to real-world applications in reef sites outside the training data. Furthermore, the larger, but sparsely annotated datasets [54, 12, 34] are essentially crowd-sourced from various annotation teams, leading to a large and noisy label set, that is challenging to transfer to new biogeographic regions. Lastly, except for the SUIM dataset [40], which provides dense segmentation labels focusing on robot-human interaction on diving missions and has no dedicated classes for live corals, no existing segmentation datasets are general-purpose in the sense of diverse camera angles, distances to the substrate, environmental conditions (water turbidity, color, lighting), and with all relevant classes visible on the images annotated.
@@ 2.2 Coral Reef Classification & Segmentation
3 Dataset
3.1 Data Collection
All imagery was collected during scuba dives at 35 sites in coral reefs of Djibouti, Eritrea, Sudan, Jordan, and Israel using GoPro Hero 10 cameras. The annotated images are video frames originally taken at 1080 $\times$ 1920px resolution and 30 frames per second in the linear lens setting. The videos were taken in a diverse set of reefs: from thriving reefs very high live coral cover to devastated environments with severe bleaching and coral mortality, from high visibility in good lighting conditions to turbid scenes in low-light settings. As the videos were acquired during transect surveying campaigns of the Transnational Red Sea Center hosted at EPFL, Coralscapes includes images including transect tools, transect lines, divers, and some human-made structures commonly encountered in reef monitoring settings. Additional information about the data collection (location, depth, reef types) is found in Appendix A. To protect the coral reefs in Coralscapes from overtourism or illegal poaching activities, and in accordance with local research permit authorities, image sites are withheld and instead replaced by a site ID.
3.2 Annotation
@@
To align to all those requirements, the annotation was done in a speculation-free manner, enforcing a conservative semantic segmentation in terms of taxonomic depth. For example, when a branching coral was too far away from the camera to decide with certainty whether it is alive or dead, it was labeled as ‘background’. Viewed closer in another video frame, the same coral may be assigned to its growth form (branching, massive, etc.) and whether it is alive or dead. A precise genus label was assigned only when the genus was clearly distinguishable. In-depth details of the annotated classes is given in Appendix A.1.
Figure 2: Histogram of the image area covered by annotations (top) and of the number of polygons present per image (bottom), highlighting the complexity of semantic segmentation in Coralscapes, compared to the CoralSCOP dataset [91]. Further comparison to the CoralSCOP dataset can be found in Appendix C.
@@
Annotation of the selected 39 classes with segmentation masks was done by four experts on coral reefs who were also present during data collection, and six additional trained annotators. The specific image frames for annotation were selected from over 200 hours of video material in order to capture the diversity of the scenes, while obtaining a good coverage of the 39 classes. The CVAT annotation interface [25] was used, which allowed annotators to interact with the Segment Anything Model [47]. While this helped to segment some objects with salient boundaries, it proved to be largely unhelpful for classes with less apparent boundaries and those appearing as small polygons, where polygons where drawn by hand. The annotation process emphasized high-quality annotations, meaning that there was a focus on annotating each frame with many polygons and on covering much of the image area, instead of annotating many images sparsely, as shown in Fig. 2. All polygons were visually verified to be correctly delineated and assigned the correct class.
In terms of class presence, the annotations of Coralscapes capture the strong imbalance in reef scenes: despite making an effort to prioritize the selection of frames with rare classes, the most common class has over 400 times more annotated polygons than the rarest class, as shown in Figure 1. Similarly, the granularity and size of the visible benthic classes varies dramatically: even though a class may appear very often, its total number of annotated pixels may still be low because it often appears in small polygons, and vice versa, as highlighted in Figure 3. For example, the median size of a ‘fish’ polygon is more than 120 times smaller than the median ‘background’ polygon. This imbalance in terms of representation and scale contributes to making Coralscapes a challenging benchmark for semantic segmentation. Additional statistics of Coralscapes (pixel counts, class frequency at image level) are provided in Appendix A.
@@ 3.3 Dataset Structure
[…]
4 Benchmarking
@@
Table 2: Quantitative results on Coralscapes segmentation, best shown in bold, second best underlined.
@@
SegFormer - MiT-b2
SegFormer - MiT-b5
Figure 4: Qualitative samples from the Coralscapes test set. Additional samples provided in Appendix D.
@@
Table 3: Transfer learning on UCSD Mosaics using 5 / 10 / 25 / 300 / all labeled pixels per image. Models marked with (Ours) are pre-trained on Coralscapes.
@@
Whenever available, we follow the original authors’ training procedures on Cityscapes in terms of hyperparameters, augmentations, and inference/evaluation strategy, unless otherwise specified. Implementation details for all benchmarked models are provided in Appendix B. To choose the number of training epochs, we train each model on the training set only and evaluate on the validation set. We then re-train each model on the train+validation sets with the number of epochs leading to the highest validation mIoU (smoothed across 10 evaluation epochs in order to reduce the effect of noise). Quantitative results are shown in Table 2, where we report performance of the models averaged over last three evaluated epochs, in order to reduce noise without training an ensemble of classifiers or cherry-picking results.
@@
While the SegFormer models showed the best performance, DINOv2 models lead to competitive pixel accuracy despite lower mIoU. This likely stems from the fact that the DINOv2 models predict the segmentation maps for the image at once, where the global structure of the image helps determine the slightly subjective delineation of ‘background’, as opposed to the strided patch predictions from SegFormer, which processes the image at higher resolution and can predict the fine-grained classes better. We find that deviating from the training procedure of SegFormer on the Cityscapes dataset by increasing the strength of data augmentations, the performance can be further improved, yielding the best model.
5 Transfer Learning & Applications
UCSD Mosaics Transfer Learning
@@ 5 Transfer Learning & Applications
We evaluate models trained on Coralscapes in a transfer learning setting on the UCSD Mosaics dataset [28], which consists of 16 orthomosaics densely annotated with 33 benthic classes and one class for the remaining substrate. We follow [4, 70, 71] in dividing the orthomosaics into patches of size $\times512$px, and remove patches with corrupted ground-truth masks [70], leaving 3974 training and 696 test patches. We evaluate both in a dense segmentation setting, as well as in a commonly performed sparse-to-dense setting, in which the training annotations are sparsified, but evaluation is performed on the unchanged dense labels. In existing works [4, 70, 71], this setting is tackled by a two-step procedure, where in a first step, algorithms propagate the sparse labels into denser masks, and in a second step, a segmentation neural network is trained on the obtained denser masks. Our approach is conceptually simpler: we train segmentation models directly on the sparsified labels, ignoring the unlabeled pixels. We select 5, 10, 25, and 300 pixel locations per orthomosaic patch uniformly at random, comparing models pre-trained on Coralscapes (with the last layer re-initialized) against off-the-shelf models with a pre-trained backbone. The results, shown in Table 3, show that pre-training on Coralscapes substantially improves segmentation with DeepLabV3+, particularly in highly sparse label regimes (5, 10, 25). When fine-tuning all parameters of a SegFormer MiT-b2, pre-training on Coralscapes improves performance in the sparse regimes, but not in the less sparse or dense (“All” in Table 3) settings. When training only the lightweight decoder on the MiT-b5 encoder weights obtained by pre-training on Coralscapes, we observe a consistent performance increase compared to an ImageNet pre-trained encoder. Interestingly, even an off-the-shelf DeepLabV3+ trained on sparse randomly located labels outperforms existing two-step procedures in sparse label regimes, despite [71] showing that selecting pixels on a grid or with a human in the loop instead of randomly is beneficial.
Crown of Thorns Starfish Survey
@@
Table 4: COTS detection results in a leave-one-out test-set setting for each of the three videos.
@@
Acanthaster planci, colloquially known as the Crown of Thorns starfish (COTS), is a common reef inhabitant in the entire Indo-Pacific, where it preys on hard corals. While in normal population sizes, COTS are a healthy part of the ecosystem, there are outbreaks of COTS, in which their population increases dramatically and the hard coral cover declines, which can lead to devastating consequences for entire reefs [36, 69]. Interventions against COTS outbreaks are most effective in the early stages of an outbreak [68, 85]: computer vision can help to rapidly detect such outbreaks by automatically analyzing imagery from reefs for COTS abundance. In [53], a dataset of three videos from manta tows (a snorkeler with a camera, towed behind a boat) in Australia were annotated with bounding boxes of COTS. The videos contain 12347, 11374, and 10759 frames at 720 $\times$ 1280 px with 3065, 6384, and 2449 bounding-box annotations of COTS respectively. We use the bounding-boxes as segmentation labels to fine-tune a SegFormer-MiT-b5 pre-trained on Coralscapes for 10 epochs, with bounding boxes being drawn around connected components of the prediction for evaluation. The results (Table 4) show that pre-training on Coralscapes leads to a substantial improvement compared to using an ImageNet pre-trained backbone. Compared to a Yolo-V8-L baseline, the transfer-learned segmentation model consistently improves the mAP@50, but the mAP@50-95, which is more sensitive to the bounding box locations, is lower. This is likely due to the naive approach to transforming segmentations into bounding boxes.
Underwater 3D Mapping from Videos
Even though 3D photogrammetry is widely used to map coral reefs [17, 49, 80, 73, 16, 7, 15], commonly used pipelines are designed for in-air mapping and often struggle in underwater environments. The detrimental effects of the water column include changes in the color of objects relative to the distance to the camera and induced blur [3, 76]. Dynamic objects such as moving fish or divers also lead to artifacts in the 3D reconstructions or even complete failure. Therefore, image collections are usually highly curated [6] to exclude images with such objects or with the water column visible.
6 Conclusion and Future Work